OpenStudy (anonymous):

find dy/dx: sin(mx) over x I need suggestions mcosx^2(mx)-(x-1)/(X^2)^2 NOT SO SURE

OpenStudy (raden):

loos still wrong

OpenStudy (raden):

*looks

OpenStudy (amistre64):

consider sin(3x) * x^(-1) as a product rule

OpenStudy (raden):

use the division rule, also the chain rule f = u/v then f' = (vu' - uv')/v^2

OpenStudy (raden):

hint : y = sin(mx)/ x u = sin(mx) ----> u' = mcos(mx) v = x ----> v' = 1 now distribute all into formula above

OpenStudy (anonymous):

I can see The bottom one is wrong